How would you solve the problem below using Cramer's Rule:
x+3y=-4
y+x=0

Answer by nyc_function(2741)

How would you solve the problem below using Cramer's Rule:
x+3y=-4
y+x=0
First switch x and y for y + x = 0.
x + 3y = -4
x + y = 0
We need the discriminant D, the discriminant of x written Dx and the discriminant of y written Dy.
D = 1 x 1 - 1 x 3
D = 1 - 3
D = -2
=============================
We now find Dx.

Dx = -4 x 1 - 0 x 3
Dx = -4 - 0
Dx = -4
=============================
We now find Dy.
Dy = -4 x 1 - 1 x 0
Dy = -4 - 0
Dy = -4
=============================
We now divided Dx and Dy by D individually.
Dx/D = x coordinate.
Dy/D = y coordinate.

x = -4/-2
x = 2.
y = -4/-2
y = 2
The solution to this system of linear equations in two variables is the point
(2, 2).
At that point, both equations cross each other.